# Polarization fringes in optical systems: a compendium

### v1

@@
Polarization fringes in optical systems: a compendium
@@
Spectral and spatial fringes in polarized light are produced by the interference of transmitted and reflected waves at the interface between materials with different indexes of refraction. These instrumental artifacts can affect the accuracy of optical designs conceived for high-sensitivity spectroscopy and polarimetry. We review the fundamental mechanisms that are responsible for these artifacts and the possible design pathways that allow us to mitigate them. In order to do so, we also present an approximate treatment of the problem of the transmission and reflection of light through (possibly absorptive) birefringent layers, relying on a few fundamental results that can be found in the already extensive literature on the subject. Unfortunately, many of these results remain the domain of a niche of investigators working in the field of thin films and optical coatings, and are often overlooked even by experienced designers of spectro-polarimetric instrumentation. The treatment presented in this work is limited to isotropic materials and uniaxial crystals, which are the most common types of optics employed in polarimetric instrumentation, and it fundamentally relies on the approximation of small birefringence for its implementation. An extensive set of modeling examples is provided to highlight the salient characteristics of polarization fringes, as well as to assess how approximate treatments such as this compare to exact but more computational expensive formulations of the problem such as Berreman’s calculus.
@@ 1 Introduction
Figure 1: The propagation of a beam through a stack of plane-parallel uniaxial crystals, rotated at different angles $\alpha$ with respect to the plane of incidence (PoI) $\langle x,z\rangle$ (in this example, $\alpha_{m}$ is negative, while $\alpha_{m+1}$ is positive). The xyzxyz axes provide the natural reference for the polarization decomposition of the incoming beam, with $y\parallel\bm{s}$, and $\bm{p}\in\langle x,z\rangle$. The figure shows the case of a stack of uniaxial crystals that have all been cut such that the optic axes ($\bm{e}$) are parallel to the interfaces. This is a common requirement for birefringent optics employed in polarimetric instrumentation, as it maximizes the birefringence experienced by the beam while minimizing polarization aberrations due to the phenomenon of double refraction. For a positive uniaxial crystal, the $\bm{o}$ and $\bm{e}$ axes in the figure correspond, respectively, to the fast and slow axes of the crystal, and they also correspond to orthogonal states of linear polarization for a ray propagating along $z$; for a negative uniaxial crystal, $\bm{o}$ becomes the slow axis and $\bm{e}$ the fast axis. The figure also clarifies the convention followed in this work, where the angle $\alpha$ of the optic is defined by the orientation of the $\bm{o}$ axis. In our treatment, we will always assume the approximation of small birefringence [8], so the ordinary (o) and extraordinary (e) rays, into which an incoming ray splits at the interface with a birefringent medium, are assumed to always lie on the PoI, and with polarizations that are orthogonal to each other, regardless of the angle $\phi$ of incidence and the orientation $\alpha$ of the optic.
2 General formalism: Jones calculus approach
@@
where we indicated with $(\cdot)^{\ast}$ the operation of complex conjugation. Next, we observe that it is possible to write
@@ 2.1 Anisotropic and rotated optics
In general, we want to be able to treat a stack of birefringent materials that are oriented at different angles $\alpha_{m}$ from the PoI (see Fig. 1), which requires the definition of the proper rotation operators necessary to model the transfer across interfaces. In our treatment, we will always assume that a birefringent optic is a perfeclty plane-parallel (i.e., with no wedge) uniaxial crystal, which has been cut such that the interface between the material and the adjacent medium is either parallel to the optic axis (called “A-cut” crystals; see description of Fig. 1) or perpendicular to it (called “C-cut” crystals).
@@
and the matrices $\mathbf{\Lambda}_{m}$ and $\mathbf{O}_{m-1,m}$ have the same structures as in Eq. (9), with all quantities referenced to the principal-axes system of the $m^{\textrm{th}}$ optic. Since the Fresnel coefficients in the $\mathbf{O}_{m-1,m}$ matrix contain the indexes of refraction of both adjacent materials, the indexes of refraction of the preceding $(m-1)^{\textrm{th}}$ optic must be referenced (or “resolved”, using the terminology of [9]) to the reference frame of the $m^{\textrm{th}}$ optic, when computing $\mathbf{O}_{m-1,m}$. In our formalism, resolving the refractive indexes of the $(m-1)^{\textrm{th}}$ material to the principal axes of the $m^{\textrm{th}}$ material is an approximate operation, which is responsible for the most significant deviations from an exact treatment of the problem [4, 3]. The downside is that particular care must be taken in the numerical implementation of our formalism, as described in App. B; the upside is a much higher computational efficiency of the approximate algorithm compared to exact treatments such as Berreman’s calculus, in order to obtain sufficiently accurate results for fringe estimation in polarimetric instruments.
@@
As we remarked at the beginning of this section, the former equations have been derived for the case of normal incidence to the stack of optics. When the incoming beam has an incidence angle $\phi_{m}>0$ to the surface normal, the rotation angle $\alpha_{m}$ in the equations above must be replaced by the “projected” rotation angle $\psi_{m}$ seen by the ray, which is given by (cf. [8], Eqs. (69))
@@
Figures 2 and 3 show examples of various optical and polarization properties of different stacks of birefringent optics, modeled with the formalism presented above. The results of Fig. 2 can be directly compared with those presented in [10, 11].
Figure 4: Wavelength dependence between 200 and 1000 nm of the (intensity normalized) Mueller matrix of a PCM optimized for full-Stokes polarimetry between 400 and 1000 nm. The polarization fringes are calculated with a spectral resolution of 20000. The PCM design uses three compound retarders in the configuration MgF2-SiO2-MgF2, where the two MgF2 elements are identical. The gray curves plotted over the fringes represent the ideal Mueller matrix from the PCM design.
A further validation of the formalism is provided by the comparison of the wavelength dependent Mueller matrix derived from it with its ideal behavior when neglecting internal wave interference and Fresnel losses. The ideal Mueller matrix for a given birefringent-stack design can be computed using the techniques illustrated by [12]. Figure 4 shows the example of a PCM consisting of a MgF2-SiO2-MgF2 stack, and optimized for full-Stokes polarimetry between 400 and 1000 nm. The ideal Mueller matrix of this design is represented by the gray curves. We note how the ideal-design solution perfectly overlaps with the fringe patterns calculated with our formalism (with the obvious exception of the transmittance element $II$, which is affected by Fresnel losses that have not been normalized in this plot).
@@ 2.2 The effect of finite f/#
Figure 5: Geometric constructs to determine the elemental solid angle (left) and phase delay (right) for a ray incident at an angle $\phi$ on the stack of optics of radius $R$, assuming a spherical wavefront with $\mathrm{f/\#}=2\tan\phi_{R}$.
As pointed out in the previous section, the formalism we propose can also be applied to the case of a ray pencil with an incidence angle $\phi$ with the surface normal of the stack different from zero. This allows us to extend the application of this formalism to the case of shaped beams, in particular, for a converging or diverging spherical wavefront produced by an ideally stigmatic optical system, in order to model the spatial-dependent fringes (such as Haidinger’s fringes; see, e.g., [7], §5.17) that are produced on an image plane with such beam configurations. Estimation of fringe amplitudes in real astronomical instruments must account for averaging over these spatial fringes [13, 14], empirically verified in [15]
@@
The elemental surface highlighted in the left panel of Fig. 5 is given by
$\delta S(\phi,\beta)$ $=$ $\frac{(r_{\phi}+\delta r_{\phi})+r_{\phi}}{2}\,\delta r_{\phi}\,\delta\beta=\frac{1}{2}\left[(r_{\phi}+\delta r_{\phi})^{2}-r_{\phi}^{2}\right]\delta\beta$
$=$ $\frac{\delta\beta}{2}\,\frac{R^{2}}{\tan^{2}\phi_{R}}\left[\tan^{2}(\phi+\delta\phi)-\tan^{2}\phi\right]\;.$
The elemental solid angle is then proportional to
$\delta\Omega(\phi,\beta)=\frac{\delta S(\phi,\beta)}{\pi R^{2}}$ $=$ $\frac{\delta\beta}{2\pi}\,\frac{\tan^{2}(\phi+\delta\phi)-\tan^{2}\phi}{\tan^{2}\phi_{R}}$ (24)
$=$ $\frac{\delta\beta}{\pi/2}\,\frac{\tan^{2}(\phi+\delta\phi)-\tan^{2}\phi}{(\mathrm{f/\#})^{2}}\;,$
where in the second row we noted that the denominator can also be expressed in terms of the f/# of the beam (see caption of Fig. 5).
The phase delay of the ray incident at $\phi$ with respect to the ray at normal incidence is $\psi_{\phi}=2\pi\,\xi/\lambda$, where $\lambda$ is the radiation wavelength, and the segment $\xi$ is highlighted in the right panel of Fig. 5. From that geometry, we easily derive $\xi=R\,(\sec\phi-1)/\tan\phi_{R}$, hence
$\psi_{\phi}=2\pi\,\frac{R}{\lambda}\,\frac{\sec\phi-1}{\tan\phi_{R}}=4\pi\,\frac{R}{\lambda}\,\frac{\sec\phi-1}{\mathrm{f/\#}}\;.$ (25)
@@ 2.3 Absorptive materials
3 The impact of polarization fringes on optical-design trades
To illustrate how the ability to model polarization fringes in layered materials can inform the design choices for polarimetric instrumentation, we show explicitly how the choices of material thicknesses (including those of non birefringent media) and beam shapes can be used to mitigate the impact of fringes on polarimetric accuracy. In particular, we analyze the behavior of polychromatic modulators (PCM [12]) optimized to work between some prescribed spectral ranges with a required spectral resolution.
@@ 3 The impact of polarization fringes on optical-design trades
The examples that follow are meant to illustrate these various trades.
3.1 A PCM for the far-UV
@@
3.2 Fringe mitigation by use of glass windows
@@
When allowed by the spectral requirements of the instrument, the choice of bias thickness of the birefringent elements, compatibly with their fabricability, and of the thickness of the possible gaps between elements (whether these are filled with index matching materials or not), can effectively be used to control the impact of intensity and polarization fringes in the instrument. Similarly, one can often utilize glass windows of adequate thickness to achieve the same results. The final example we present demonstrates such utilization of windows made of isotropic materials for the purpose of decreasing the dominant fringe period well below the spectral resolution of the instrument. For this example, we consider the same SiO2-MgF2 achromatic half-wave retarder used for Fig. 7, illuminated by a collimated beam in normal incidence around 500 nm. The three different curves in the Mueller matrix plots of Fig. 13 show the cases of the original achromat model (black curves), and for the same achromat sandwiched between two index-matching glass windows (nominally, amorphous SiO2 and MgF2 crystals) with a thickness of 5 mm (orange curves) and 15 mm (blue curves). The sampling resolution of the plot is 20000, for an effective spectral resolution of about 10000. We note the significant suppression of polarization fringes that can be attained by utilizing windows of adequate thickness in applications that do not require a very high spectral resolution.

### v2

# On the modeling and mitigation of interference fringes in polarimetric instrumentation

@@
On the modeling and mitigation of interference fringes in polarimetric instrumentation
@@
Spectral and spatial fringes in polarized light are produced by the interference of transmitted and reflected waves at the interface between materials with different indexes of refraction. These instrumental artifacts can affect the accuracy of optical designs conceived for high-sensitivity spectroscopy and polarimetry. We consider the principal sources of these artifacts and the possible design pathways to mitigate them. In order to do so, we have developed an approximate yet agile treatment of the problem of the transmission and reflection of light in birefringent materials, which fundamentally relies on the assumption of small birefringence of the modeled materials for its implementation. The comparison of our results with those from more rigorous treatments, such as Berreman calculus, thus also serves as a test of the limits of the small-birefringence approximation in optical design applications. The treatment presented in this work is limited to isotropic materials and uniaxial crystals, which are the most common types of optics employed in polarimetric instrumentation. An extensive set of modeling examples is provided to illustrate the salient characteristics of polarization fringes and their dependence on optical design parameters.
@@ 1 Introduction
Figure 1: Propagation of a light ray through a stack of plane-parallel uniaxial crystals, rotated at different angles $\alpha$ with respect to the plane of incidence (PoI) $\langle x,z\rangle$ (in this example, $\alpha_{m}$ is negative, while $\alpha_{m+1}$ is positive). The xyzxyz axes provide the natural reference for the polarization decomposition of the incoming beam, with $y\parallel\bm{s}$, and $\bm{p}\in\langle x,z\rangle$. The figure shows the case of a stack of uniaxial crystals that have all been cut such that the optic axes ($\bm{e}$) are parallel to the interfaces. This is a common requirement for birefringent optics employed in polarimetric instrumentation, as it maximizes the birefringence experienced by the beam while minimizing polarization aberrations due to the phenomenon of double refraction. For a positive uniaxial crystal, the $\bm{o}$ and $\bm{e}$ axes in the figure correspond, respectively, to the fast and slow axes of the crystal, and they also correspond to orthogonal states of linear polarization for a ray propagating along $z$; for a negative uniaxial crystal, $\bm{o}$ becomes the slow axis and $\bm{e}$ the fast axis. The figure also clarifies the convention followed in this work, where the angle $\alpha$ of the optic is defined by the orientation of the $\bm{o}$ axis. In our treatment, we will always assume the approximation of small birefringence [8], so the ordinary (o) and extraordinary (e) rays, into which an incoming ray splits at the interface with a birefringent medium, are assumed to always lie on the PoI, and with polarizations that are orthogonal to each other, regardless of the angle $\phi$ of incidence and the orientation $\alpha$ of the optic.
2 Jones calculus approach to polarization fringes
@@
where we indicated with $(\cdot)^{\ast}$ the operation of complex conjugation. Next, we observe that
@@ 2.1 Anisotropic and rotated optics
In general, we want to be able to treat a stack of birefringent materials that are oriented at arbitrary angles $\alpha_{m}$ from the PoI (see Fig. 1), which requires the definition of the proper rotation operators necessary to model the transfer across interfaces. In our treatment, we will always assume that a birefringent optic is a perfeclty plane-parallel (i.e., without wedge) uniaxial crystal, which has been cut such that the interface between the material and the adjacent medium is either parallel to the optic axis (called “A-cut” crystals; see description of Fig. 1) or perpendicular to it (called “C-cut” crystals).
@@
and the matrices $\mathbf{\Lambda}_{m}$ and $\mathbf{O}_{m-1,m}$ have the same structures as in Eq. (9), with all quantities referenced to the principal-axes system of the $m^{\textrm{th}}$ optic. Since the Fresnel coefficients in the $\mathbf{O}_{m-1,m}$ matrix contain the indexes of refraction of both adjacent materials, the indexes of refraction of the preceding $(m-1)^{\textrm{th}}$ optic must be “referenced” (or “resolved”, using the terminology of [9]) to the principal-axes frame of the $m^{\textrm{th}}$ optic, when computing $\mathbf{O}_{m-1,m}$. Such a resolution of the refractive indexes of the $(m-1)^{\textrm{th}}$ material to the principal axes of the $m^{\textrm{th}}$ material is an approximate operation of our formalism, which is responsible for the most significant deviations from an exact treatment of the problem [3, 4]. The downside is that particular care must be taken in the numerical implementation of the formalism, as described in App. B; the upside is a demonstrated higher computational efficiency of the approximate approach, when compared with some numerical implementations of the exact theory such as Berreman calculus, while it still provides sufficiently accurate results for quantitative fringe estimation in polarimetric instruments.
@@
As we remarked at the beginning of this section, the former equations have been derived for the case of normal incidence on the stack of optics. When the incoming beam has an incidence angle $\phi_{m}>0$ to the surface normal, the rotation angle $\alpha_{m}$ in the equations above must be replaced by the “projected” rotation angle $\psi_{m}$ seen by the ray, which is given by (cf. [8], Eqs. (69))
@@
Figure 4: Wavelength dependence between 200 and 1000 nm of the (intensity normalized) Mueller matrix of a polychromatic modulator (PCM) optimized for full-Stokes polarimetry between 400 and 1000 nm. The polarization fringes are calculated with a spectral resolution of 20000. The PCM design uses three compound retarders in the configuration MgF2-SiO2-MgF2, where the two MgF2 elements are identical. The gray curves plotted over the fringes represent the ideal Mueller matrix from the PCM design.
@@
Figures 2 and 3 show examples of various optical and polarization properties of different stacks of birefringent optics, modeled with the formalism presented above. In particular, the results of Fig. 2 can be directly compared with those presented in [10, 11].
A further validation of the formalism is provided by comparing the predicted wavelength-dependent Mueller matrix with the one obtained in the ideal case when wave interference and Fresnel losses are neglected. Figure 4 shows the example of a “polychromatic” modulator (PCM [12]) consisting of a MgF2-SiO2-MgF2 6-plate stack, and optimized for full-Stokes polarimetry between 400 and 1000 nm. The ideal Mueller matrix of this design is represented by the gray curves. We note how the ideal solution perfectly overlaps with the fringe patterns calculated with our formalism, with the obvious exception of the transmittance element $II$, which is affected by Fresnel losses that have not been normalized in this plot.
All these figures demonstrate the close connection among intensity, retardance, and polarizance/diattenuation fringes. This is clearly manifested at specific conditions of constructive and destructive interference of the transmitted and reflected waves, which is attained for $m\pi/2$ values of the retardance, with $m$ an integer. In particular, Fig. 2 shows how retardance and polarizance/diattenuation fringes are suppressed around half-wave conditions, while intensity fringes, in both transmission and reflection, tend to be suppressed around quarter-wave conditions. This is also clearly demonstrated by Fig. 3, where the occurrence of nulls and maxima of the fringe amplitudes in intensity and polarizance/diattenuation across the spectrum appear to be in “opposition of phase”, with the two nulls of polarizance fringes specifically occurring at the two wavelengths where the compound retarder is exactly half-wave.
@@ 2.2 The effect of finite f/#
Figure 5: Geometric constructs to determine the elemental surface (left) and phase delay (right) pertaining to a ray incident with an angle $\phi$ on the stack of optics of radius $R$, assuming a spherical wavefront with $\mathrm{f/\#}=(2\tan\phi_{R})^{-1}$.
As pointed out in the previous section, the formalism we propose can also be applied to the case of a ray pencil with an incidence angle $\phi$ with the surface normal of the stack different from zero. This allows us to extend the application of this formalism to the case of shaped beams, in particular, for a converging or diverging spherical wavefront produced by an ideally stigmatic optical system, in order to model the spatial-dependent fringes (such as Haidinger’s fringes; see, e.g., [7], §5.17) that are produced on an image plane with such beam configurations. Estimation of fringe amplitudes in real astronomical instruments must account for averaging over these spatial fringes [13, 14], as empirically verified in [15].
@@
The elemental surface shown in the left panel of Fig. 5 is given by
$\delta S(\phi,\beta)$ $=$ $\pi\Bigl[r_{\phi}^{2}-(r_{\phi}-\delta r_{\phi})^{2}\Bigr]\frac{\delta\beta}{2\pi}=\pi R^{2}\,\frac{r_{\phi}^{2}-(r_{\phi}-\delta r_{\phi})^{2}}{R^{2}}\,\frac{\delta\beta}{2\pi}$
$=$ $\pi R^{2}\,\frac{\tan^{2}\phi-\tan^{2}(\phi-\delta\phi)}{\tan^{2}\phi_{R}}\,\frac{\delta\beta}{2\pi}\;,$
and thus the corresponding elemental solid angle is
$\delta\Omega(\phi,\beta)=\frac{\delta S(\phi,\beta)}{\pi R^{2}}=\frac{\tan^{2}\phi-\tan^{2}(\phi-\delta\phi)}{\tan^{2}\phi_{R}}\,\frac{\delta\beta}{2\pi}\;.$ (24)
The phase delay of the ray incident at $\phi$ with respect to the ray at normal incidence is $\psi_{\phi}=2\pi\,\xi/\lambda$, where $\lambda$ is the wavelength, and the segment $\xi$ is highlighted in the right panel of Fig. 5. From that geometry, we easily derive $\xi=R\,(\sec\phi-1)/\tan\phi_{R}$, hence
$\psi_{\phi}=2\pi\,\frac{R}{\lambda}\,\frac{\sec\phi-1}{\tan\phi_{R}}\;.$ (25)
@@
It is important to remark that the effect of the phase delay $\psi_{\phi}$ cancels out when calculating spatially resolved Mueller-matrix maps of the type shown in Figs. 6 and 7, and therefore it can simply be neglected for such a purpose.
@@ 2.3 Absorptive materials
To conclude this section, we want to comment on the sensitivity of fringes to the extinction coefficient $k$ of absorptive materials. The values of $k$ range by orders of magnitude across wavelength, they typically are affected by larger experimental errors than in the case of the real index $n$, especially for moderately absorptive materials, and they also can vary quite widely between different material grades. At the end of Sect. 2.1 we commented on the interdependence of intensity and polarization fringes, and how this manifests particularly around specific values of the retardance. Thus, it can be expected that both intensity and polarization fringes will generally be impacted by the absorption characteristics of a material, and how uncertainties in the value of $k$ may impact the modeling accuracy of interference fringes is an important question to address. Our finding is that, by measuring optical transmission with the typical accuracy of commonly available tools, as in the case of the DKIST polarizer-window example of Fig. 9, we are able to constrain the actual $k$ value precisely enough for fringe modeling purposes. In summary, when absorption values are large enough to impact fringe characteristics, the corresponding $k$ values are sufficiently well constrained by standard optical transmission measurements for accurately modeling both intensity and polarization fringes. On the other hand, for cases with very small and often highly uncertain $k$ values, the impact of absorptivity on fringes is normally negligible.
3 The impact of polarization fringes on optical-design trades
Using the modeling framework derived in the previous section, we now want to illustrate how the ability to model polarization fringes in layered materials can inform the design choices for polarimetric instrumentation. In particular, we want to show explicitly how the choices of material thicknesses (including those of any non birefringent media in the stack) and beam shapes can be used to mitigate the impact of fringes on polarimetric accuracy. In order to do so, we analyze the behavior of PCMs optimized to work between some prescribed spectral ranges with a required spectral resolution.
@@ 3 The impact of polarization fringes on optical-design trades
The examples presented in the next section are meant to illustrate these various trades. A practical note about these modeling examples is necessary, which concerns the specification of the polar grid to adopt for optimally sampling the fringe pattern across the beam. Spatial-fringe patterns such as those of Figs. 68 demonstrate that polarization fringes may become rapidly variable along the angle-of-incidence $\phi$ of the ray within the beam, whereas the $\beta$-azimuth variations around the surface normal typically have a much slower frequency, usually with periods that are multiple of $\pi/2$, as it can be expected for homogeneous A-cut and C-cut birefringent crystals. This suggests that the sampling of the $\beta$-domain can typically be significantly coarser than the sampling of the $\phi$-domain, in order to properly capture the fringe variations along that coordinate, thus allowing for the optimization of the computational effort. For the examples illustrated in the next section, fringe averages over the beam extent appeared to be stable for $\delta\phi\lesssim 0.025^{\circ}$ and $\delta\beta\lesssim 5^{\circ}$. Obviously, in the case of less symmetric patterns, such as those that are produced in the presence of an overall tilt of the beam entering the optic, may need significantly finer $\beta$ samplings, and so the choice of the sampling grid must be assessed case by case.
@@
3.1 A PCM for the far-UV
Figure 13: Same modeling of interference fringes as in Fig. 11, assuming a f/13 beam configuration and a spectral sampling resolution of 40000, using the same PCM model in Table 1 (blue curves; identical solution as the blue curves in Fig. 11), and after reducing the bias thickness of the four waveplates to 0.3 mm (orange curves).
@@ 3.1 A PCM for the far-UV
3.2 Fringe mitigation by material thickness adjustments
@@
When allowed by the spectral requirements of the instrument, the choice of bias thickness of the birefringent elements, compatibly with their fabricability, and of the thickness of the possible gaps between elements (whether these are filled with index matching materials or not), can effectively be used to control the impact of intensity and polarization fringes in the instrument. Similarly, one can often utilize glass windows of adequate thickness to achieve the same results. The final two examples we present demonstrate the trades of choosing adequate bias thicknesses of birefringent elements and the utilization of windows made of isotropic materials, for the purpose of decreasing the dominant fringe period well below the spectral resolution of the instrument.
For the first example, we consider again the FUV PCM modeled in the previous section, and evaluate again the amount of fringe suppression for a f/13 beam of radiation spectrally sampled with an instrument resolution of 40000—similarly to what was shown in Fig. 11—after reducing the bias thickness of the PCM plates from 0.4 mm down to 0.3 mm (cf. Table 1). The blue curves in Figs. 11 and 13 correspond to the same solution, whereas the orange curves in Fig. 13 now correspond to the case of reduced bias thickness. This comparison shows that, for the adopted shape of the incoming beam and the specified spectral resolution of the instrument, the adoption of thinner plates for the same PCM model leads to a significantly worse suppression of the fringe pattern over the wavelength interval of this example.
For the second example, we consider again the SiO2-MgF2 achromatic half-wave retarder used for Fig. 7, illuminated by a collimated beam in normal incidence around 500 nm. The three different curves in the Mueller matrix plots of Fig. 14 show the cases of the original achromat model (black curves), and for the same achromat sandwiched between two index-matching glass windows (nominally, amorphous SiO2 and MgF2 crystals) with a thickness of 5 mm (orange curves) and 15 mm (blue curves). The sampling resolution of the plot is 20000, for an effective spectral resolution of about 10000. We note the significant suppression of polarization fringes that can be attained by utilizing passive windows of adequate thickness in applications that do not require a very high spectral resolution.